### v1

# Understanding Software Engineering Agents: A Study of Thought-Action-Result Trajectories

Abstract
Large Language Model (LLM)-based agents are increasingly employed to automate complex software engineering tasks such as program repair and issue resolution. These agents operate by autonomously generating natural language thoughts, invoking external tools, and iteratively refining their solutions. Despite their widespread adoption, the internal decision-making processes of these agents remain largely unexplored, limiting our understanding of their operational dynamics and failure modes. In this paper, we present a large-scale empirical study of the thought-action-result trajectories of three state-of-the-art LLM-based agents: RepairAgent, AutoCodeRover, and OpenHands. We unify their interaction logs into a common format, capturing 120 trajectories and 2822 LLM interactions focused on program repair and issue resolution. Our study combines quantitative analyses of structural properties, action patterns, and token usage with qualitative assessments of reasoning coherence and feedback integration. We identify key trajectory characteristics such as iteration counts and token consumption, recurring action sequences, and the semantic coherence linking thoughts, actions, and their results. Our findings reveal behavioral motifs and anti-patterns that distinguish successful from failed executions, providing actionable insights for improving agent design, including prompting strategies, failure diagnosis, and anti-pattern detection. We release our dataset and annotation framework to support further research on transparent and robust autonomous software engineering agents.
I Introduction
LLMs are becoming increasingly popular to automate various software engineering tasks. Successful applications of LLMs include code completion [1, 2, 3, 4], automated program repair [5, 6], test case generation [7, 8, 9], test oracle generation [10, 11], and fuzz testing [12]. These approaches typically operate by prompting an LLM with a query, either a single time or in a hard-coded feedback loop.
Beyond querying LLMs with fixed prompt templates and within hard-coded algorithms, LLM agents have recently emerged as another promising approach to software engineering automation. Such agents are autonomous systems that use LLMs to iteratively reason about a problem, invoke existing tools, and adapt based on outputs produced by these tools [13]. For example, LLM agents can autonomously fix bugs and other issues [14, 15, 16, 17], write tests to reproduce issues reported in natural language [18, 19, 20], and automate complex project setups [21, 22]. To address a given task, an LLM agent autonomously plans and executes a sequence of actions, leveraging its reasoning capabilities and external tools. Agents in software engineering typically use tools such as code editors, compilers, debuggers, and test runners to interact with the code base and the environment.
While the success of LLM agents in software engineering tasks is promising, the underlying decision-making processes remain largely opaque. However, as their adoption increases, a critical question emerges: How do LLM agents reach their solutions, and why do they (not) work well on specific tasks? Answering this question is crucial for understanding the strengths and limitations of LLM agents, as well as for improving their performance and reliability. In particular, insights into the usage of tools could help to improve the design of LLM agents and their integration with existing software engineering tools. Likewise, insights into the reasoning processes of agents could reveal common pitfalls and best practices, which will provide a basis for improving future LLM agents.
Understanding how LLM agents address software engineering tasks is important but also challenging. On the upside, LLM agents provide detailed logs that document their reasoning and the actions they take. We call these logs agent trajectories, which we define as a sequence of thought-action-result triples, where each triple consists of a reasoning step, or thought, expressed by an LLM, an external action (e.g., invoking a compiler), and the result of that action (e.g., the compiler output). While the availability of agent trajectories is a promising starting point, there currently is no established methodology to systematically study them.
This paper presents the first in-depth empirical study of thought-action-result trajectories of LLM-based software engineering agents. We focus on agents for automated program repair and issue solving, as this problem has received significant attention recently, and study trajectories from three state-of-the-art agents, namely RepairAgent [14], AutoCodeRover [15], and OpenHands [16]. Our study addresses three research questions:
  • RQ1: What are the overall properties of trajectories, such as the number of iterations and consumed LLM-tokens?
  • RQ2: What kinds of actions do agents perform and what sequential patterns emerge within trajectories?
  • RQ3: How do agents develop and adjust their reasoning and actions throughout task execution, and how coherent are thoughts, actions, and feedback?
As a cross-cutting theme, we also investigate how the answers to the above questions differ for trajectories that successfully complete a task versus those that fail.
To address these questions, we introduce a novel methodology that analyzes LLM-based software engineering agents through their thought-action-result trajectories. We begin by collecting logs from the three studied agents (RepairAgent, AutoCodeRover, OpenHands) and convert them into a unified trajectory format. We then compute statistics to examine trajectory properties and how to relate to success or failure of an agent (RQ1). Next, we manually categorize agent actions, and mine frequent action sequences (RQ2). Finally, to study reasoning coherence (RQ3), we apply open coding to label semantic relationships between thoughts, actions, and results. This mixed-method approach enables us to extract both quantitative insights and qualitative interpretations of agent behavior. Our study leads to several key findings:
  • Test-driven repair agents (RepairAgent, OpenHands) have longer, token-intensive failing trajectories, reflecting task complexity; conversely, AutoCodeRover’s streamlined retrieve-locate-fix workflow is more efficient.
  • Successful action sequences balance exploration, explanation, fix generation, and testing; failures exhibit repetitive, non-adaptive cycles and debugging anti-patterns like repeated identical actions without follow-up or fix generation without intermediate testing.
  • Semantic alignment between thoughts and actions is essential; even rare misalignments strongly correlate with failure or increased computational cost, underscoring the need for explicit verification mechanisms (e.g., self-reflection or critique) to ensure coherence.
  • Agents differ in reasoning dynamics: flexible state-machine agents enable revisiting and steady progress; test-driven agents prioritize systematic verification; streamlined agents focus on concise retrieve-and-fix workflows. Identifying the optimal architecture for each task warrants further research.
Our work relates to prior efforts in LLM interpretability, log analysis, and trajectory-based agent improvement. LLM interpretability techniques provide insights into why a model produces a specific output [23, 24], e.g., via counterfactual reasoning [25] or by highlighting specific code locations [26]. Unlike these approaches, our work focuses on the trajectory of an agent’s reasoning and actions over time, rather than just a single prediction. To help users inspect agent behavior, prior work has proposed tools for visualizing the interactions within multi-agent systems [27, 28]. Our work complements these efforts by focusing on a single agent that interacts with its environment, and by targeting agents in software engineering. Finally, several techniques aim at fully automatically improving agents via reinforcement learning based on past trajectories [29, 30, 31]. Instead, our work provides insights that enable agent developers to understand and improve agents.
In summary, this paper contributes the following:
  1. 1.
    Methodology. We introduce a novel trajectory-based modeling and analysis capturing the interplay of thoughts, actions, and results in SE agents.
  2. 2.
    Empirical results. We provide the first systematic study of agent trajectories across multiple state-of-the-art SE agents, revealing characteristic patterns and challenges.
  3. 3.
    Insights and implications. We offer insights into agent decision-making processes that inform better prompting strategies, trajectory supervision, and evaluation metrics.
II Methodology
Figure 1: Overview of our methodology.
To answer the research questions given in Section I, we present a novel methodology for analyzing the trajectories of LLM agents for software engineering. Figure 1 provides an overview of our methodology, which consists of five main components. Given the raw logs of an LLM agent, we first parse the logs into a unified representation of trajectories. We then address RQ1 by computing statistics, e.g., about the length of trajectories and the costs imposed by the agents. To study the actions performed by the agents (RQ2), we systematically inspect and categorize the actions, followed by mining the resulting action sequences for sequential patterns. Finally, we analyze the semantic relationships between thoughts, actions, and result (RQ3) through open coding, allowing us to study the coherence and consistency of an agent’s behavior. The remainder of this section describes our data collection and each of the components of our methodology in detail.
II-A Data Collection
We collect trajectory logs from three SOTA SE agents:
  • RepairAgent [^1] https://github.com/sola-st/RepairAgent [14]: An autonomous agent for program repair evaluated on 835 bugs of Defects4J [32].
  • AutoCodeRover [^2] https://github.com/AutoCodeRoverSG/auto-code-rover [15]: An agent designed for automated issue resolution, evaluated on the 300 instances of SWE-bench Lite [33]. We utilize the trajectories from AutoCodeRover V1. Although subsequent versions of AutoCodeRover exist, their trajectories were unavailable to us. Additionally, AutoCodeRover conducts three runs and reports average results; since there is no fundamental difference among these runs, we selected the trajectories from the first run for this study.
  • OpenHands [^3] https://github.com/All-Hands-AI/OpenHands [16]: A general-purpose code editing agent, also evaluated on the 300 tasks in SWE-bench Lite [33]. We use trajectories obtained from OpenHands CodeAct, a version that is tailored for issue resolution.
To ensure a diverse yet manageable dataset for detailed manual annotation, we randomly sample 40 trajectories per agent. Given the average success rate of approximately 27% across the three agents, we proportionally sample about 10 successful trajectories per agent. Each trajectory captures a complete bug repair or issue resolution through iterative agent-environment interactions. Our dataset thus comprises 120 trajectories totaling 2822 iterations across all agents.
II-B Trajectory Parsing and Representation
To ensure consistency in our analysis, all selected trajectories are processed using our trajectory parser. which converts the raw logs from different agents into a unified representation:
Definition 1.
An agent trajectory is a sequence of iterations that each consist of three components:
$\mathcal{T}=[(t_{1},a_{1},r_{1}),(t_{2},a_{2},r_{2}),\dots,(t_{n},a_{n},r_{n})]$
where each tuple $(t_{i},a_{i},r_{i})$ represents the following:
  • The thought $t_{i}$ is a natural language description of the agent’s internal reasoning, such as a diagnostic insight, an explanation of program behavior, or a proposed fix.
  • The action $a_{i}$ is an invocation of an external operation, such as editing code, running a test, or another tool.
  • The result $r_{i}$ is the response to action $a_{i}$, such as compiler output, test results, execution traces, or error logs.
To convert raw logs into a structured format, we implement parsers for each agent. RepairAgent and OpenHands produce structured JSON logs with clearly separated reasoning, actions, and results, each requiring tailored parsing due to distinct JSON schemas. For AutoCodeRover, we develop a heuristic to map its semi-structured output into trajectories by detecting tool calls via tool names, extracting preceding reasoning as thoughts, and capturing the call results as separate text items.
II-C Statistical Analysis
To address RQ1, we compute several trajectory-level metrics that capture the overall properties of agent trajectories. Given a trajectory $\mathcal{T}=[(t_{1},a_{1},r_{1}),\dots,(t_{n},a_{n},r_{n})]$, we compute the following metrics.
Definition 2.
The trajectory length of $\mathcal{T}$ is the number $|\mathcal{T}|$ of iterations in the trajectory.
The trajectory length reflects task complexity, with easier tasks completed in fewer iterations. To manage computational effort, agents limit trajectory length, e.g., RepairAgent caps at 40 iterations, while OpenHands allows up to 100.
Definition 3.
The trajectory cost of $\mathcal{T}$ is the total number of tokens consumed by the agent during the trajectory:
[…]
The trajectory cost is important because queries to LLMs are typically charged based on the number of tokens processed.
Definition 4.
The success of a trajectory $\mathcal{T}$ is a binary label indicating the agent’s success in completing the task.
We define success as in reported in the original evaluation of the agents, i.e., whether the agent produced a valid patch or solution to the task. Specifically, for AutoCodeRover and OpenHands, we consider a trajectory successful if the agent produces a patch that passes the tests executed by SWE-bench [33]. For RepairAgent, we consider a trajectory successful if the produced patch passes the tests executed by Defects4J [32] (called “plausible”) and manual inspection confirms that the patch is equal or semantically to the developer-produced patch (called “correct”).
II-D Categorizing Actions
To address RQ2, which investigates the actions performed by agents, we categorize the actions in the trajectories. A key challenge is the diversity of actions performed by different agents, which includes various operations, such as code editing and executing tests, while not following a unified naming convention to refer to actions. After an initial inspection of trajectories, we observe that the actions performed by agents can be grouped into a small number of high-level categories that reflect activities developers commonly perform when debugging or fixing code [34].
Based on this observation, we categorize each action into one of the following eight categories:
  1. 1.
    Explore: Listing, navigating, and reading files to gather contextual information.
  2. 2.
    Locate: Identifying specific code regions or elements relevant to the bug or issue.
  3. 3.
    Search: Employing search tools to find occurrences of code elements, symbols, or patterns expressed via regular expressions.
  4. 4.
    Reproduce: Creating test cases to confirm the presence of the bug or the issue.
  5. 5.
    Generate fix: Proposing or writing code modifications intended to resolve the issue.
  6. 6.
    Run tests: Executing test cases to validate the proposed fix or to validate the test cases themselves.
  7. 7.
    Refactor: Improving code quality through optimization, reformatting, or commenting.
  8. 8.
    Explain: Producing natural language explanations or summaries of the issue or applied fix.
To categorize the actions, we first map known tools used by agents such as search tools into the respective category (e.g, Search). Then, we manually inspect the remaining actions that we were unable to map automatically (e.g, a terminal command). The labeling is initially performed by one author, followed by a discussion of ambiguous and uncertain cases to resolve them collaboratively among multiple authors. A small subset of actions (8.3%) does not fit into any of the above categories, such as internal state transitions in RepairAgent and LLM-suggested actions that AutoCodeRover fails to parse. We exclude these cases from subsequent analyses.
II-E Mining Sequential Action Patterns
We analyze sequences of categorized actions in agent trajectories to identify recurrent decision-making patterns. Each trajectory is represented as a sequence of action labels, e.g., Explore $\rightarrow$ Locate $\rightarrow$ Generate Fix $\rightarrow$ Reproduce $\rightarrow$ Run Tests. We study these sequences by mining for frequently occurring subsequences. First, we implement a lightweight Python script to extract fixed‑length $n$‑grams and compute their frequencies, choosing $n$ large enough to avoid trivial short sequences but small enough to retain sufficient diversity and easy-to-analyze patterns; from our experiments, $n\in\{4,5,6\}$ performs best, and for this study we set $n=4$, which yields the richest set of recurring patterns. Next, for each agent we identify the top‑frequency 4‑grams and analyze how their relative frequencies differ between successful and failed tasks. Finally, we flag patterns overrepresented in failing trajectories as well as rare or unusual 4‑grams that may indicate nonstandard or suboptimal workflows.
II-F Labeling Semantic Relationships
TABLE I: Semantic relationships between parts of agent trajectories.
Relationship Definition and example
Thought $t_{i}$rightarrow action $t_{i}$ (within one iteration):
Alignment The action logically aligns with the thought. Example: Thought “I need to inspect usages of function foo” aligns with action “search_calls(’foo’)”.
Misalignment The action does not reflect the thought. Example: Thought “I need to write a fix for this bug” misaligns with action “read documentation”.
Thought $t_{i}$rightarrow thought $t_{i+1}$ (across consecutive iterations):
Follow-up The thought is a logical, intuitive follow-up to the previous one. Example: “Line 161 indicates a potential off-by-one error” followed by “I should verify and adjust the loop boundary at line 161”.
Refinement The thought adds specific details to the previous thought. Example: “I should focus on collecting information to fix the bug based on the expected and actual outputs in the failing test case” refines “I should focus on the assertion at line 342 in ExtensionFunctionTest.java involving comparing the expected and actual output values”.
Redundancy The thought is repeated without added new insights. We count both semantic and syntactic repetition. Example: “The bug is in function A” is redundant with “The bug seems to happen in function A.”
Divergence The thought shifts focus to a different direction. Example: “I will now suggest a fix…” diverges from “I need to find the location of the bug.”
Contradiction The thought contradicts the previous thought. Example: “I think the error is in function A” contradicts “The error is actually in function B, as shown by the stack trace”.
Action $a_{i}$rightarrow action $a_{i+1}$ (across consecutive iterations):
Follow-up The action is a plausible follow-up to the previous action. Example: Running test cases after writing a fix.
Refinement The action improves or extends the previous action. Example: “Search for ’error handling”’ is refined to “Search for ’error handling in function X”’.
Repetition The same action is taken again without modification with the same parameters (syntactically or semantically). Example: Running the same test case multiple times without changing the code.
Divergence The action does not follow naturally from the previous one. Example: “Trying a candidate patch” followed by “Making search queries”.
Result $r_{i}$rightarrow thought $t_{i+1}$ (across consecutive iterations):
Follow-up The thought is a logical follow-up on the results of the previous iteration. Example: Result “Test passed” followed by thought “Since test cases passed, I think I can end my task”.
Refinement The result provides specific details that refine the next thought. Example: Detailed compiler errors refine the hypothesis about the bug location.
No influence The result does not affect the next thought. Example: Result “New search results: …” followed by thought “I need to suggest a fix”.
Misinterpretation The thought misinterprets the previous result. Example: An error output followed by thought “The bug was fixed”.
Result $r_{i}$rightarrow action $a_{i+1}$ (across consecutive iterations):
Informative The result provides information that is incorporated into the next action. Example: A specific error message is used as an argument in the next code search.
Triggering The result directly triggers a particular action. Example: Successful bug fixing triggers the end of task.
No influence The result has little or no effect on the next action. Example: The LLM obtains new search results but none of them affect the fix suggested next.
RQ3 investigates the semantic relationships among trajectory components (thoughts, actions, results) to evaluate the coherence and consistency of an agent’s behavior. We systematically label relationships both within iterations (e.g., thought to action) and across iterations (e.g., how a result in iteration $i$ influences the thought in iteration $i+1$).
To label relationships, we apply an open coding approach [35] to the trajectories. Open coding is a qualitative analysis technique where a researcher identifies and labels recurring patterns within a dataset. Beginning with no predefined labels, the coding evolves through manual inspection as new patterns emerge. We apply this technique to five types of relationships between components of agent trajectories. The result of this process is the set of relationships presented in Table I. The following summarizes the five types of relationships we consider, and examples of the relationships we identify.
Thought $t_{i}$$\rightarrow$ action $a_{i}$
These relationships capture the degree to which an action that the agent choses faithfully implements the immediately preceding thought expressed by the agent. For example, we assign an Alignment label when the action logically follows from the thought, such as thinking “I need to inspect usages of foo” followed by searching the code base for occurrences of foo.
Thought $t_{i}$$\rightarrow$ thought $t_{i+1}$
These relationships describe how the agent’s internal reasoning evolves across iterations. For example, we label a pair of thoughts as Follow-up when the second thought is a natural continuation of the first, such as “I need to inspect usages of foo” followed by “I should check the function definitions that use foo”. In contrast, we label a pair of thoughts as Divergence when the second thought pivots to a different line of reasoning, such as “I need to inspect usages of foo” followed by “I will now suggest a fix for the bug”.
Action $a_{i}$$\rightarrow$ action $a_{i+1}$
These relationships characterize how the agent’s operations progress toward task completion. For example, we label a pair of actions as Repetition when the agent performs the same action consecutively without new effect on the agent and codebase.
Result $r_{i}$$\rightarrow$ thought $t_{i+1}$
These relationships characterize how the agent incorporates feedback from its environment into the next thought. For example, we label a pair of result and thought as Follow-up when the thought is a logical continuation of the previous result, such as “Tests passed” followed by “Since test cases passed, I think I can end my task”. In contrast, we assign the label Misinterpretation when the agent understands a result incorrectly, such as “Test failed” followed by “The bug was fixed”.
Result $r_{i}$$\rightarrow$ action $a_{i+1}$
These relationships capture how the agent’s results influence its subsequent actions. For example, we label a pair of result and action as Informative when the result provides information that is incorporated into the next action (e.g, api name found through search).
We focus on these five specific types of relationships because they directly capture the immediate and most influential semantic connections within and across consecutive iterations. Long-term sequence relationships or more distant dependencies can be informative but not straightforward to systematically label and interpret due to their extension and complexity. Our targeted scope balances depth and feasibility.
The relationships in Table I are the result of labeling 40 trajectories from each of the three agents, which totals to 14K pairs of trajectory components. Initially, some of the authors label each pair, while discussing ambiguous and uncertain cases with other authors to resolve them collaboratively.
III Results
This section presents a detailed analysis of the three research questions introduced earlier.
III-A RQ1: Trajectory-Level Properties
We investigate trajectory properties aiming to understand how their iterative processes differ in length and complexity. We answer this by quantitatively analyzing the distribution of iteration counts and tokens consumption.
Figure 2: Comparative trajectory length between successful and unsuccessful trajectories (y-axis differs between the agents).
Figure 2 shows the distribution of iteration counts across agents. RepairAgent has a mean trajectory length of 34 iterations, indicating consistent iterative reasoning. Similarly, OpenHands averages 29 iterations, reflecting comparable depth. In contrast, AutoCodeRover (ACR) completes workflows in fewer steps (6 on average), highlighting a more efficient patch generation process.
The figure also compares trajectory lengths between successful and unsuccessful instances, revealing agent-specific patterns. For RepairAgent, unsuccessful cases have longer trajectories (mean 40) than successful ones (mean 22), indicating higher effort on difficult tasks, consistent with its design to continue attempts until a fix is found or early termination is triggered. OpenHands shows a similar trend, though a bit less, with some unsuccessful trajectories reaching the 100-iteration cap. For AutoCodeRover, successful and unsuccessful trajectories have similar average lengths, but unsuccessful ones exhibit higher variability and outliers, often due to internal errors like Response Parse Error. AutoCodeRover terminates after patch application without integrating tests into its workflow.
Discrepancies in Trajectory Length Test-driven repair leads to longer trajectories, especially in unsuccessful cases, reflecting sustained exploration for solutions. In contrast, agents with direct patch application show shorter trajectories, with unsuccessful cases marked mainly by greater variance.
Figure 3: Input, Output, and Total tokens over successful trajectories and unsuccessful ones (y-axis is different for each agent).
Figure 3 presents cumulative token consumption per trajectory. AutoCodeRover exhibits the smallest footprint (mean 23K, median 14K tokens). OpenHands averages about 1.2M tokens (median 900K), approximately 52 times more than AutoCodeRover. RepairAgent demonstrates moderate usage (mean 220K, median 180K). Key observations are as follows:
  • RepairAgent: Unsuccessful debugging attempts use more tokens than successful ones, reflecting extended iterations and additional context (e.g., search).
  • OpenHands: Successful trajectories consume more input tokens than unsuccessful ones, while output token usage is similar (more input context used in successful runs).
  • AutoCodeRover (ACR): Trajectories are consistently short and efficient; failures often involve Error Response Parsing sequences that modestly increase token usage.
Token Consumption Patterns RepairAgent’s token usage increases with trajectory length due to its iterative nature. OpenHands’s higher token usage in successful runs suggests that input context aids success. AutoCodeRover maintains efficient consumption, with variability mainly from parsing errors.
These findings indicate that both iteration count and total token consumption provide complementary and insightful metrics for assessing problem complexity and the behavioral dynamics of each agent. Subsequent research questions dive deeper into the precise characterization of agents behavior.
III-B RQ2: Actions and Patterns of Actions Sequences
III-B1 Overall Action Usage
Figure 4(a) compares the overall action categories distribution. Across all agents, the most frequent actions are Generate Fix (23%), Run tests (20%), Explain (20%), and Explore (18%). While these core actions are common, each agent exhibits distinct patterns in action selection and sequencing.
RepairAgent tends to emphasize a balanced approach incorporating search and exploration, fix generation, and comprehensive testing. In contrast, OpenHands places particular emphasis on a test-driven methodology, including the generation of issue-reproducing tests, which is distinctive to OpenHands’s workflow. AutoCodeRover operates through a structured search–locate–fix process and typically omitting explicit Reproduce and Test steps in its main pipeline. It is worth noting that a subsequent extension, SpecRover [36], introduces dedicated reproduction and testing phases.
III-B2 Action Usage Over Task Progress
To analyze behavior evolution, we normalize iterations $i\in[1..N]$ of a trajectory to percentage progress. Figures 4(b)4(d) depict the proportion of actions per category at each progress point (e.g., 20%). In RepairAgent (Fig. 4(b)), the initial 0–20% of the trajectory is evenly split between Fix Generation plus Running Tests ( 50%) and the remaining categories (Explore, Explain, Search, Locate). From 20% to 100%, Fix Generation and Running Tests dominate ( 70%). OpenHands (Fig. 4(c)) starts with  50% Explore and Search in the first 20%, then shifts towards increased Fix Generation, Reproduce, and Running Tests, which constitute  75% in the last 20%. AutoCodeRover (Fig. 4(d)), with median $N=5$, spends the first 50% on Search, Locate, and Explain, then transitions to Locate (30%) and Generate Fix (20%) in the latter half. Failing trajectories include many ErrorResponse events, omitted from the plot.
Takeaways from Agent Actions Developers can use actions distribution as a secondary metric in evaluating the quality of agents by analyzing the frequency and timestamps of actions or by comparing them to real trajectories obtained from human experts.
(a) Actions frequency across agents.
(b) RepairAgent
(c) OpenHands
(d) AutoCodeRover
Figure 4: Normalized action distribution across task execution phases for each agent.
III-B3 Sequence Patterns
To complement the time-distributed action plots, Figures 5, 6, 7 presents a visualization of the top most frequent 4-gram sequences executed by each agent, with each row representing a distinct sequence and its corresponding frequency in successful and unsuccessful trajectories.
In the case of RepairAgent, successful 4-grams prominently feature cycles that alternate between Generate Fix and exploratory actions such as Explain or Explore, with little dominance by any single pattern, as shown in Figure 5. In contrast, unsuccessful trajectories often show a strong concentration in repetitive Generate Fix cycles as shown in the last row of Figure 5 (e.g., multiple consecutive Generate Fix actions).
For OpenHands (Figure 6), the most frequent successful sequences balance between context gathering, patch generating, and testing (generate and run). Unsuccessful trajectories, on the other hand, show more repetitive cycles of exploration and search actions (e.g, first rows in the figure), or sequences of repeated testing and fixing steps without interleaving actions. This highlights the importance balancing between different actions and having strategic shift in debugging phases instead of continuous repetition of same actions.
For AutoCodeRover (Figure 7), the most common 4-grams in successful trajectories typically involve sequences with Explain and Explore, often followed by Generate Fix. Notably, successful sequences tend to cluster around varied exploratory and explanatory actions.
Figure 5: Top 11 sequences of actions 4-grams (RepairAgent).
Figure 6: Top 12 sequences of actions 4-grams (OpenHands).
Figure 7: Top 11 sequences of actions 4-grams (AutoCodeRover).
Actions Sequences Analysis of frequent 4-gram action sequences shows that successful trajectories balance exploration, explanation, and validation steps, while unsuccessful ones exhibit repetitive, non-adaptive action cycles.
In addition to sequence patterns linked to success, we identified debugging anti-patterns, i.e, action sequences causing incoherent results, resource waste, or unclear objectives, akin to trajectory smells described in [37]. Examples include:
1) Repeated identical actions without follow-up: For instance, repeated Search actions occur across all three agents. While some repetitions arise from unsuccessful prior searches, others neglect to act on found results. RQ3 elaborates on this no-influence phenomenon.
2) Repeated “Generate Fix” without testing: Observed mainly in OpenHands and AutoCodeRover, unlike RepairAgent which mandates testing after fix generation. Testing is crucial to identify necessary patch modifications to understand if a fix works and why.
3) Task termination without proper test validation: Although AutoCodeRover omits test validation by design, OpenHands sometimes generates reproducing tests without verifying their effectiveness (i.e, the ability of the test to reveal the bug), leading to unreliable patch validation at the end.
Debugging Anti-Patterns We identified recurrent debugging anti-patterns that reduce effectiveness: (1) repeated identical actions without follow-up, causing wasted effort; (2) consecutive Generate Fix steps without intermediate testing, delaying patch evaluation; and (3) premature task termination without a proper test validation. Robust agents design should include checks for anti-patterns and trajectory smells.
III-C RQ3: Semantic Relationships
The semantic modeling in Sec. II-F offers a structured lens to dissect agent decision-making. It enables us to identify where planning fails, execution diverges, or the agent does not learn from its outputs. Here, we provide deeper analysis to highlight tendencies and differences among agents.
TABLE II: Semantic relationships between components of agent trajectories (NI = No-Influence, Ctr=Contradiction).
Agent Thought–Action Thought–Thought Action–Action Result–Thought Result–Action
Aln Mis Ctr Div Fol Red Ref Div Fol Ref Rep Fol MisInt NI Ref Inf NI Trig
RepairAgent 98.6 1.4 1.4 6.3 69.7 9.2 13.5 17.8 51.4 20.2 10.6 75.4 3.4 3.4 17.8 30.0 5.6 64.4
- Success 99.2 0.8 1.4 4.2 75.1 5.2 14.2 26.1 56.7 11.8 5.3 82.5 4.0 1.7 11.8 43.5 0.9 55.6
- Failure 98.6 1.4 1.4 6.6 68.9 9.8 13.4 16.7 50.7 21.3 11.3 74.5 3.3 3.6 18.6 27.4 6.5 66.1
OpenHands 96.8 3.2 0.1 8.1 80.2 1.6 10.1 23.8 61.2 13.7 1.2 70.6 3.3 6.8 19.3 30.3 5.1 64.6
- Success 99.0 1.0 0.0 5.9 78.6 1.8 13.6 28.6 58.7 12.2 0.5 73.4 2.8 2.8 21.0 19.6 2.7 77.8
- Failure 96.5 3.5 0.1 8.3 80.4 1.6 9.7 23.3 61.6 13.9 1.3 70.3 3.3 7.4 19.0 31.4 5.3 63.3
ACR 96.3 3.7 0.0 1.4 72.4 0.9 25.3 20.8 69.7 3.9 5.6 59.9 0.0 6.2 33.9 37.7 25.8 36.5
- Success 100.0 0.0 0.0 2.1 70.8 0.0 27.1 19.5 78.0 2.4 0.0 59.7 0.0 8.1 32.3 46.7 21.7 31.7
- Failure 95.2 4.8 0.0 1.2 72.8 1.2 24.9 21.2 67.2 4.4 7.3 59.9 0.0 5.7 34.4 34.8 27.2 38.0
Table II presents the distribution of semantic relationships across trajectories for the three agents. Columns represent key relational categories (e.g., Divergence, Follow-up, Refinement…). Rows show the prevalence of each relation for both successful and unsuccessful trajectories.
III-C1 Analysis of Thought-Action Relationships
An agent’s thought is intended to guide the subsequent action and its parameters. If the action does not align with the thought, the agent may be misled. For example, if a thought suggests running tests relevant to a bug, but the action executes unrelated tests, the agent may incorrectly conclude the issue is resolved when the tests pass while the correct tests were not executed.
Among the three agents, misalignment between thought and action is rare but closely associated with failing trajectories. For example, OpenHands shows 1% misalignment in successful and 3.5% in failing trajectories, while AutoCodeRover has 0% in successful and 4.8% in failing ones. Although not all misalignments result in failure, even a single instance can significantly increase trajectory length or cause failure. Furthermore, misalignment is more prevalent in failing trajectories: only 1 out of 10 successful RepairAgent trajectories contain a misalignment compared to 12 out of 30 failing ones.
An example of misalignment between intent and action occurs at iteration 5 during RepairAgent’s execution on bug Compress_13. The agent, after identifying a plausible fix, attempted refinement but instead suggested an empty fix. This error caused confusion and extended the process from iteration 6 (where the correct patch was available) to iteration 38, increasing computational cost. Although such misalignment is uncommon and model-dependent, incorporating mechanisms to verify that actions align with articulated thoughts is prudent. Prior work has explored this for LLMs and agents [38, 39]. Newer reasoning models (e.g., o3, deepseek, Claude 4) embed self-reflection and criticism. Integrating these into SE agents, where precise translation of natural language to actions is critical, would improve robustness and efficiency in a significant fraction of trajectories.
Thought-Action Alignment Even a single misalignment between thought and action can cause failure or increase computational cost. We recommend explicit verification alignment, e.g, through self-reflection, critique frameworks, or reasoning models.
III-C2 Analysis of Thought-Thought Relationships
Unlike chain-of-thought reasoning, LLM-agent thoughts are interleaved with actions and their outcomes, influencing subsequent thoughts. Table II shows that most thoughts are follow-ups or refinements of previous ones, though distinct patterns emerge per agent. For example, RepairAgent has the highest number of contradicting thoughts, which are rare in OpenHands and AutoCodeRover. GPT-3.5, used by RepairAgent, often makes trivial mistakes. In bug Compress_13, RepairAgent states at iteration 5: The fix was successful, and all test cases passed…, but contradicts this at iteration 6: The previous fixes did not work, and the bug is still present…. Additionally, RepairAgent frequently produces redundant thoughts, especially in failing trajectories, often repeating generic statements like The previous fix failed, I need to suggest a new one without explaining failure reasons or next steps.
Divergence between thoughts is notable in RepairAgent (37/40 trajectories) and OpenHands (34/40), with both positive and negative effects. Diverging from promising paths may mislead the agent, while diverging from unproductive loops enables exploration and balances exploitation and exploration. For example, in OpenHands’s Django-10924, iterations 1–6 focus on FilePathField implementation, but iteration 7 shifts to reproducing the issue, breaking repetitive code reading and exploring testing and fixing instead.
Relationships between consecutive thoughts Follow-ups and refinements illustrate iterative reasoning in LLM agents. Frequent contradictions and redundancies lead to confusion and resource waste. Developers should minimize contradictions and redundancies, while balancing exploitation and exploration using example trajectories or mutation-like mechanisms.
III-C3 Analysis of Consecutive Actions Pairs
Across all three agents, Follow-up relationships dominate, indicating most actions build on their immediate predecessors. However, the distribution of other relationship types varies among agents. Divergence is prominent in Action-Action relationships and is often beneficial, reflecting transitions between phases (e.g., from bug localization to fix suggestion). This is evident as action divergence is more common in successful trajectories (e.g., 26 vs. 16 for RepairAgent). Conversely, repetitive sequences and consecutive action repetition, particularly in RepairAgent and AutoCodeRover, strongly correlate with failing trajectories. Repetition wastes iterations and may induce unproductive loops, as also observed in RQ2. Repetition can be detected using sequence analysis methods such as sliding window matching or n-gram frequency analysis. Detected repetitions can then be addressed by diversification strategies that encourage exploration of alternative actions or transitions.
Actionable Insights on Action Sequence Dynamics Repetition can be mitigated using sequence analysis methods, applying diversification strategies, including alternative actions or exploration steps to break repetitive cycles and cover new paths.
III-C4 Influence of Results on Thoughts and Actions
Analysis of Result–Thought Relationships
Ideally, agents should interpret results accurately, using informative signals to guide their reasoning, correct errors, or refine strategies. However, Table II shows both strengths and limitations in how agents process and respond to results.
Across all agents, the most common relationship is follow-up, where a thought logically builds on the received result, such as responding to a failed test by suggesting a new patch. Refinement relationships are also frequent, indicating efforts to improve the solution based on results. Misinterpretations and no influence cases are less common but occur more often in failing trajectories, reflecting instances where agents ignore or misunderstand result signals. For example, in RepairAgent’s run on bug Cli_11 at iteration 6, the agent misinterprets the result indicating that the last fix was empty and instead of suggesting a non-empty fix, it moves on to analyzing the code with the intention to change empty function arguments.
Analysis of Result–Action Relationships
The outcome of an action should shape the agent’s next action. In successful runs, we observe a high proportion of triggering actions, where the result directly prompts a relevant next step. Informative actions are also common, as agents refine their actions based on information drawn from previous results.
However, failures often correlate with a higher proportion of no-influence actions (0.9 vs 6.5 for RepairAgent), where the agent’s next move does not reflect the information content of the preceding result. These observations suggest that effective LLM agents not only need to interpret results accurately at the reasoning level but also translate those insights into appropriately responsive action sequences.
IV Limitations and Threats to Validity
Our study offers insights into LLM-based agents in software engineering tasks but faces limitations, notably the potential subjectivity in our annotation and classification process despite inter-annotator agreement checks. We evaluate three bug-fixing agents which, despite their common use, do not fully capture all software development tasks or real-world deployment scenarios. The agents rely on distinct models, ensuring diversity but complicating fair comparisons. The generalization of our results to other SE agents or tasks, such as code generation or testing, remains open for study.
Furthermore, the sampled trajectories used for in-depth analysis may not capture the full variability in agent behavior. Although we apply a random sampling strategy to ensure diversity, rare but important patterns could be overlooked. Future studies could address this limitation by scaling up the dataset and using automated pattern detection techniques.
Our correlation and sequence mining analyses reveal trends and associations without establishing causality between behaviors and task success. Despite this limitation, our study provides a structured approach to analyzing LLM-based agent behavior and offers valuable insights for future research and autonomous software engineering tool development.
V Related Work
Deep Learning and LLMs in Software Engineering
Neural reasoning, particularly with LLMs, has attracted significant attention in software engineering research [40]. Early work focused on neural code completion [1, 2, 3, 4]. Recent studies extend LLMs to tasks such as program repair [5, 6, 41], issue solving [42], test case generation [7, 8, 9], test oracle generation [10, 11], and fuzz testing [12]. Unlike the agentic approaches studied in this paper, these methods prompt LLMs with pre-defined algorithms.
LLM Agents
Recently, LLM agents have emerged as a new paradigm in software engineering, where LLMs autonomously solve complex tasks by reasoning and interacting with external tools [13]. A prominent application is automated bug and issue fixing, with numerous agents proposed, including RepairAgent [14], AutoCodeRover [15], OpenHands [16], SWE-Agent [17], Magis [43], AgentCoder [44], MarsCode Agent [45], FixAgent [46], and Passerine [47]. Motivated by this task’s relevance, our work studies three representative agents. Other agents focus on writing tests to reproduce issues from natural language [18, 19, 20, 48], automating project setup [22, 49, 21], debugging computational notebooks [50], and analyzing root causes of failures from logs [51]. To support progress, several benchmarks have been introduced, such as SWE-bench [33], SWE-bench+ [52], EnvBench [53], and automated benchmark generation [54]. For a comprehensive overview, we refer to recent surveys [55, 56]. Rather than proposing a new agent, our work systematically analyzes existing agents to inform future agent design.
AI Interpretability and Explainability
Our work relates to efforts on interpreting and explaining AI models. These include techniques that identify relevant nodes and edges in graph-based neural models for specific predictions, e.g., in NLP [26, 57] or vulnerability detection [58]. Other work explains predictions via counterfactual reasoning [25] or by learning predicates that characterize mispredicted inputs [59]. Yeo et al. analyze chain-of-thought reasoning, focusing on the impact of training and feedback [60]. Surveys provide overviews of explanation techniques for black-box models [23] and LLMs [24]. Recent work also targets LLM agents, e.g., by proposing progress metrics for trajectory analysis [61], visualizing agent behavior in multi-agent systems [27], or offering debugger-like interfaces [28]. To our knowledge, our work is the first to systematically study the reasoning and decision-making processes of multiple LLM-based agents in software engineering.
Automatically Improving LLM Agents
To improve LLM agents, one approach applies reinforcement learning based on past trajectories and their outcomes [29, 30, 31]. Another direction fine-tunes LLMs using previous agent trajectories to enhance performance on specific tasks or optimize multi-agent communication [62, 63]. A third line introduces additional agents that interpret or provide feedback on the main agent’s actions [64]. While our work also aims to improve LLM agents, we focus on deriving insights through systematic analysis of existing agents and their tool interactions, rather than fully automating the improvement process.
VI Conclusion
This study systematically analyzes LLM-based autonomous agents in software engineering, focusing on decision-making, reasoning consistency, and trajectory dynamics. Using statistical analysis, semantic relation modeling, open coding, and sequence pattern mining, we identify key behavioral trends. Successful trajectories balance information gathering, hypothesis testing, and fix validation, while unsuccessful ones exhibit redundant exploration or premature fixes. Categorizing actions and sequences reveals agents’ strengths and weaknesses, guiding improvements in efficiency, reliability, and reasoning. Future work could include expanding agent and benchmark coverage, refining classification methods, and developing automatic failure detection and mitigation techniques.
Data availability
The paper is under review. We will release the data once our work is published in its final version.